An aircraft shell is fired vertically upward with a muzzle velocity of 1000ms.calculate;


a=the maximum height it can attain


b=the time taken to reach this height


c=the instantaneous velocities at the ends of 20sec and 50sec


d=when will it's height be 37.5km.interprete your result. neglect air resistance.

1
Expert's answer
2021-11-11T12:16:21-0500

(a) Let's take the upwards as the positive direction. Then, we can find the maximum height from the kinematic equation:


"v^2=v_0^2+2gh,""0=v_0^2+2gh,""h=\dfrac{-v_0^2}{2g}=-\dfrac{(1000\ \dfrac{m}{s})^2}{2\times(-9.8\ \dfrac{m}{s^2})}=51\ km."

(b) We can find the time that shell takes to reach the maximum height from the kinematic equation:


"v=v_0+gt,""0=v_0+gt,""t=\dfrac{-v_0}{g}=\dfrac{-1000\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=102\ s."

(c) We can find the instantaneous velocity at the end of "20\ s" from the kinematic equation:


"v=v_0+gt=1000\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\times20\ s=804\ \dfrac{m}{s}."

We can find the instantaneous velocity at the end of "50\ s" from the kinematic equation:


"v=v_0+gt=1000\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\times50\ s=510\ \dfrac{m}{s}."


(d) We can find the time the shell takes to reach the height of 37.5 km from the kinematic equation:


"y=v_0t-\dfrac{1}{2}gt^2,""37500=1000t-4.9t^2,""4.9t^2-1000t+37500=0."

This quadratic equation has two roots: "t_1=154.6\ s" and "t_2=49.5\ s". The time "t=49.5\ s" corresponds to the case where the shell first reaches the height of 37.5 km (as it rises up). The time "t=154.6\ s" corresponds to the case where the shell secondly reaches the height of 37.5 km (as it falls down).